### v2

@@ Abstract
Both studies reported excellent agreement with the experimental data. It was, of course, necessary to introduce units of time, by writing $t=\mathcal{W}t^{\prime}$ where $t^{\prime}$ is the dimensional time. Standley and Wright used esr to study the magnetic relaxation of a an organic copper complex (copper dipyrromethene), following a microwave pulse, fitting their data at 2.45 K (for example) with $\mathcal{W}=1.06/{\rm ms}$, $\sigma=1.57$. Roinel et al. used nmr to measure the relaxation of the polarisation (magnetization) of 169Tm nuclear spins in TmPO4 at $T=50$ mK, $H_{1}=1$ T, fitting their data with $\mathcal{W}=1/(26~{\rm hr})$ and $\sigma=17$.
@@
Figure 5: Comparison with of the theory with experimental data on a reduced timescale $t_{0}$ defined in the text. The lower figure is an enlargement of the upper one for short times. The black points are the data of Ref. [17], and the orange points are the data of Ref. [18]. The blue lines are the exact solution, Eq. 21 for (top to bottom) $\sigma=1.57$ [17], $\sigma=17$ [18] and $\sigma=500$, a large value for comparison. The magenta dashed line is the exact Fourier transform of Eq. 1.
Eq. 20 is exact and equivalent to the following exact solution for $v(t)$:
$v(t)=-\frac{1}{1+\sigma}W\left(-\frac{\sigma\left(e^{-\sigma-t}\right)^{\frac{1}{1+\sigma}}}{1+\sigma}\right)$ (21)
where $W$ is the Lambert W-function. This equation shows that curves of different $\sigma$ can be compared on a scale of reduced time $t_{0}=t/(1+\sigma)$. For the inverse Fourier transform of Eq. 1 the corresponding transformation is $t_{0}=t/\sigma$ [15], but the difference is immaterial given that the latter is a large $\sigma$ approximation.
Rescaling accordingly, the data are replotted in Fig. 5 and compared with the theory. It can be seen that the value of $\sigma=17$ (Roinel et al’s data) is already approaching the large $\sigma$ limit. This is to be expected as, following the earlier arguments, $\sigma=17$ should give over two decades of $1/\omega^{3/2}$ scaling, whereas for $\sigma=1.57$ (Standley and Wright’s data), the corresponding range would be negligible. It would be interesting to find further examples of such behavior among real systems, particularly with much larger values of $\sigma$.
@@
The value of $\alpha=3/2$ found here, equivalent to stretched exponential ($e^{-t^{\beta}}$) relaxation with $\beta=2-\alpha=1/2$, may be compared to the values found in other magnetic systems. Ising spin glasses [6] display values of $\beta\approx 0.25-0.3$. In single molecule magnets, a value of $\beta=1/2$ has been observed at millikelvin temperatures at resonant tunneling points and interpreted in terms of fluctuating random fields [8, 9]. In spin ice, $\alpha$ evolves from $2$ down to about 1.5 at subkelvin temperatures [11], where the value of roughly 1.5 has been associated with the diffusion of emergent magnetic monopoles in a spatially fractal environment [10, 20]. It is perhaps unsurprising that that the few-body systems (paramagnets, single molecule magnets) seem to exhibit exponents that are rational fractions, while the many body systems (spin glasses, spin ice) exhibit more general exponent values.
@@
It is finally worth noting some general insights afforded by the present results. Exponential relaxation indicates that the rate of some relaxing quantity is proportional to its displacement from equilibrium. In the present case $dv/dt$ is instead linear with $u$, the displacement of the phonon equilibrium (Eq. 7), but the latter is a highly nonlinear function of $v$, and this is the origin of the approximate $e^{-\sqrt{t}}$ relaxation. A situation that is mathematically (though not physically) similar arises in the case of $e^{-\sqrt{t}}$ relaxation in an ordered two-dimensional Ising model where the radius of relaxing circular domains (droplets) plays a similar role to $u$ in the phonon case [21]. In physical terms, the example of the phonon bottleneck shows how just two relaxation processes (both in themselves exponential) may generate a $1/f^{\alpha}$ power spectrum or $e^{-t^{\beta}}$ relaxation, when sufficiently strongly coupled. The anomalous $\alpha,\beta$ are associated with an emergent, non-equilibrium steady state, where a frequency-dependent linear response function can be defined. The phonon bottleneck thus illustrates how anomalous relaxation can emerge in a basic, but realistic, model of a far from equilibrium, and essentially quantum mechanical, steady state. Given its simplicity, the model might lend itself to generalisations, and given current interest in two-level systems, either as qubits [22] or as sources of noise [23], it might find contemporary relevance.

### v3

@@ IV Application to real systems
The first case is where $2|\chi|/\chi_{T}\omega$ simply represents the Fourier transform of the normalised relaxation function $v(|t|)$. If experimental relaxation data is consistent with the $F(t)$ of Eq. 13 then it is most likely consistent with its Fourier transform. There are at least two studies, by Standley and Wright [16], and by Roinel et al. [21], that tested Eq. 9 and hence Eq. 13. Both studies reported excellent agreement with the experimental data. It was, of course, necessary to introduce units of time, by writing $t=\mathcal{W}t^{\prime}$ where $t^{\prime}$ is the dimensional time. Standley and Wright used esr to study the magnetic relaxation of an organic copper complex (copper dipyrromethene), following a microwave pulse, fitting their data at 2.45 K (for example) with $\mathcal{W}=1.06/{\rm ms}$, $\sigma=1.57$. Roinel et al. used nmr to measure the relaxation of the polarisation (magnetization) of 169Tm nuclear spins in TmPO4 at $T=50$ mK, $H_{1}=1$ T, fitting their data with $\mathcal{W}=1/(26~{\rm hr})$ and $\sigma=17$.
These fits are reproduced in Fig. 4 to illustrate how closely the limiting form Eq. 13 agrees with experiment, even for small values of $\sigma$. In the Figure, curves of different $\sigma$ are compared on a scale of reduced time $t_{0}=t/(1+\sigma)$. The data of Roinel et al. ($\sigma=17$) are therefore consistent with a function whose Fourier transform closely approximates $\omega^{-3/2}$ behaviour over more than two decades.
@@
Figure 4: Replotted experimental data [16, 21] on a reduced timescale $t_{0}$ defined in the text. The lower figure is an enlargement of the upper one for short times. The black points are the data of Ref. [16], and the orange points are the data of Ref. [21]. The blue lines are the exact solution, Eq. 13 for (top to bottom) $\sigma=1.57$ [16], $\sigma=17$ [21] and $\sigma=500$, a large value for comparison. The figure illustrates how closely experimental systems approach the exact limiting form, Eq. 13.
The second case to consider is the application of a magnetic field that oscillates around zero, $H=he^{i\omega t}$. The question can be posed, can the complex susceptibility predicted here actually be measured by ac-magnetometry?
In order to address this question, it is necessary to ‘unpack’ the Faughnan-Strandberg equations somewhat. Clearly, the mathematical analysis assumes that both $\sigma$ and $\tau_{\rm sl}$ (the latter of which which sets the units of time in the equations) remain constant as $M^{0}$ varies. Ostensibly, this is only possible if the field, and hence the energy spacing of the two levels (measured by frequency $\nu\propto h$) and the parameters $\sigma$ and $\rho$, remain fixed to a good approximation. This, of course, applies to a step function field applied, but contradicts a field oscillating around zero. Nevetherless, the following relations can be established:
$\tau_{\rm sl}=\frac{1}{2\rho A(p^{0}+1/2)}=-\frac{n^{0}}{2\rho AN},$
$\sigma=-\frac{n^{0}A\tau_{\rm ph}}{\Delta\nu}.$
So, for the equations to have meaning in the case of an oscillating applied field, we must have $\rho A\propto n^{0}\propto\Delta\nu/A\tau_{\rm ph}$. Suppose that $\rho$ is fixed to its total, or average value, during the field cycle, $\rho\rightarrow\bar{\rho}$ and, for consistency with this, $\Delta\nu/\nu={\rm constant}$. Then, given that $n^{0}\propto h$ and the average $\nu\propto h$, it follows that $A\propto\nu\propto 1/\tau_{\rm ph}$. These relations, i.e. lifetimes increasing at low energy, do not seem unreasonable. Hence, while it remains far from proved, it does seem plausible that an ac-susceptibility experiment could reveal the $1/\omega^{3/2}$ behaviour predicted here.
The third, and final, case to consider is when a small oscillating field is applied to a system at equilibrium in a relatively strong applied field. In this case, the term $M/M^{0}$ multiplying $\sigma$ in Eq. 15 can be set to unity, and the relaxation is exponential [24].
V Discussion and Conclusions
To expand upon the remarks in the Introduction, an isolated system will always tend to exhibit non-exponential decay as energy and information are retained in the system, leading to a greater likelihood of ‘memory’. Conversely, connection to a heat bath, will generally suppress memory as energy and information are transferred to the bath and effectively lost to the system, leading to exponential relaxation. The phonon bottleneck, considered here, demonstrates this latter point rather transparently. If the phonons remain at thermal equilibrium then the phonon population does not depend on the history of the relaxation because it is fixed to its equilibrium value. Hence there is no ‘feedback’ to the magnetization and the decay is exponential. If, on the other hand, the phonon occupation number differs from its equilibrium value, then its decay towards equilibrium depends on the history of the decay of the spin populations: this information feeds back to affect the decay of those populations, leading to non-exponential relaxation.
@@ V Discussion and Conclusions
The value of $\alpha=3/2$ found here, loosely equivalent to stretched exponential ($e^{-t^{\beta}}$) relaxation with $\beta=2-\alpha=1/2$, may be compared to the values found in other magnetic systems. Ising spin glasses [6] display values of $\beta\approx 0.25-0.3$. In single molecule magnets, a value of $\beta=1/2$ has been observed at millikelvin temperatures at resonant tunneling points and interpreted in terms of fluctuating random fields [8, 9]. In spin ice, $\alpha$ evolves from $2$ down to about 1.5 at subkelvin temperatures [11], where the value of roughly 1.5 has been associated with the diffusion of emergent magnetic monopoles in a spatially fractal environment [10, 25]. It is perhaps unsurprising that that the few-body systems (paramagnets, single molecule magnets) seem to exhibit exponents that are rational fractions, while the many body systems (spin glasses, spin ice) exhibit more general exponent values.
Among non-magnetic systems, there is much interest in the related problem of noise in two-level systems. One example is flux noise in SQUIDS [26]. Another is the fluorescence noise of quantum dots where values of $\alpha$ and high-frequency crossovers rather similar to those found in the present work have been observed [27, 28]. In such systems, the anomalous noise originates in the connection of the two-level system to other energy states, either extrinsic [26] or intrinsic [27, 28], creating a certain analogy with the phonon bottleneck system. In this context, it would be interesting to study the magnetic noise directly in the case of the phonon bottleneck.
@@
It is finally worth noting some general insights afforded by the present results. Exponential relaxation indicates that the rate of some relaxing quantity is proportional to its displacement from equilibrium. In the present case $dv/dt$ is instead linear with $u$, the displacement of the phonon equilibrium (Eq. 8), but the latter is a highly nonlinear function of $v$, and this is the origin of the anomalous relaxation. A situation that is mathematically (though not physically) similar arises in the case of $e^{-\sqrt{t}}$ relaxation in an ordered two-dimensional Ising model where the radius of relaxing circular domains (droplets) plays a similar role to $u$ in the phonon case [29]. In physical terms, the example of the phonon bottleneck shows how just two relaxation processes, both in themselves exponential, may generate a $1/f^{\alpha}$ power spectrum relaxation, when sufficiently strongly coupled. The phonon bottleneck thus illustrates how anomalous relaxation can emerge in a basic, but realistic, model of a far from equilibrium, and essentially quantum mechanical, steady state. Given its simplicity, the model might lend itself to generalisations, and given current interest in two-level systems, either as qubits [30] or as sources of noise [31], it might find contemporary relevance.